3. The patient is a 44 lb. child who is 5½ years old. Ordered: Midazolam 1.5 mg IV On hand: Midazolam 2 mg/2mL Package insert states: For children the usual dose ranges from 0.05 -0.4 mg/kg

A. Is the dosage ordered safe? YES_____ NO_____ Show calculations that proves your choice.

B. How will the medication be given?___________

C. How often will the medication be given?___NA__________

D. How much in mL of the medication to be given at prescribed time?________________

E. What do you do if the dose is not a safe dose?_________________________________
Answer #1

ANSWER.

1) Calculation of safe dose for the patient.

       Ordered dose = 1.5 mg.

     Recommended dose = 0.05 mg - 0.4 mg/kg.

   Weight of the patient = 44 lb.

      Convert lb to kg for calculation

               1 lb = 0.45 kg.

     Weight of the patient weight 44 lb = 44 * 0.45

                                                           = 19.8 kg.

      Safe dose for a patient weight 19.8 kg = 0.05 * 19.8 - 0.4 * 19.8

                                                              = 0.99 mg to 7.92 mg.

           So

    * Ordered dose(1.5 mg) is safe for the child.

B) This medication is given slowly through intravenous route.

C) How often the medication be given = NA.

D) Calculation of amount(ml) of drug required.

           Formula = Desired dose /Dose in hand * Quantity.

     Here

         Desired dose(D) = 1.5 mg.

         Dose in hand (H) = 2mg.

            Quantity (Q) = 2 ml.

As per above formula

     Amount(ml) of drug required = D/H * Q

                                                    = 1.5 / 2 * 2

                                                    = 0.75 * 2

                                                    = 1.5 ml.

E) If the dose is not a safe dose withhold the medication and inform physician to change

     the dose.